OpenStudy (anonymous):

Find the value of y so that the line passing through (2, 6) and (1, y) has a slope of 5.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{6-y}{2-1}=5 }\]can you solve for y?

OpenStudy (anonymous):



OpenStudy (solomonzelman):

for y? no.

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{6-y}{2-1}=5 }\] \[\huge\color{blue}{ \frac{6-y}{1}=5 }\] \[\huge\color{blue}{ 6-y=5 }\]what about now?

OpenStudy (anonymous):

still not getting it

OpenStudy (solomonzelman):

Then you don't seem to have enough knowledge for this question, I am really sorry to say this, but if you need a good math tutor, bot someone like me.

OpenStudy (anonymous):

do u know the answer

OpenStudy (solomonzelman):

Me? Guess?

OpenStudy (anonymous):

what you do is add the y on the left side of the equal sign and what you do to one side you do to the other than on the right side you minus the 5 and you will get 1 = y